OpenStudy (freckles):

\[\text{ Let } m \text{ and } n \text{ be natural numbers such that } \\ A=\frac{(m+3)^n+1}{3m} \text{ is an integer } \\ \text{ Show } A \text{ is odd.} \]

OpenStudy (usukidoll):

well definition of odd is 2k+1

OpenStudy (freckles):

Assume A is even then A=2i where i is an integer. \[2i=\frac{(m+3)^n+1}{3m} \\ \text{ let's see if we can arrive at some contradiction } \\ 2i(3m)=(m+3)^n+1 \\ 2(3im)=(m+3)^n+1 \text{ so } (m+3)^n+1 \text{ is even } \\ 2(3im)-1 \text{ is odd } \text{ so } (m+3)^n \text{ is odd } \] thinking a bit more... (not sure if I'm going to get this to work)

OpenStudy (freckles):

having trouble with contradiction ...was thinking about writing it like this but I don't see how to use vieta's jumping of whatever it is called if I can even use that ) \[A =\frac{(m+3)^n+1}{3m} \in \mathbb{Z} \\ A(3m)=(m+3)^n+1 \\ 0=(m+3)^n+1-A(3m) \\ (m+3)^n+1-3Am=0\]

ganeshie8 (ganeshie8):

The given problem is equivalent to : \[\dfrac{m^n+3^n+1}{3m} =A \in \mathbb{Z} \implies \text{A is odd}\]

OpenStudy (freckles):

you did mod 3

OpenStudy (freckles):

why does doing that given us an equivalent problem

ganeshie8 (ganeshie8):

\[(m+3)^n = m^n+3^n + 3m(\text{stuff})\]

ganeshie8 (ganeshie8):

Hmm yeah not exactly same as it disturbs the parity of \(A\)

OpenStudy (freckles):

actually found an answer but I don't like it because I don't understand it :p

OpenStudy (freckles):

do you want me to share it?

ganeshie8 (ganeshie8):

yeah im almost stuck haha

OpenStudy (freckles):

click "This LTE thing is pretty cool"

OpenStudy (freckles):

where I have looked up LTE and I believe it just means lifting the exponent lemma/theorem thingy

OpenStudy (freckles):

if m is even why does that allow us to do mod 3?

OpenStudy (freckles):

By the way I got my problem from here if you wanted to know : https://www.google.com/url?sa=t&rct=j&q=&esrc=s&source=web&cd=20&cad=rja&uact=8&ved=0CFQQFjAJOAo&url=http%3A%2F%2Fwww.math.muni.cz%2F~bulik%2Fvyuka%2Fpen-20070711.pdf&ei=suSIVZ-ZL4G9sAWMh7QY&usg=AFQjCNESVFKertjKo1RRDenoVvOsgsul0w (by the way do you know if there is a way to hide how long this link actually is using some kind of LaTeX voodoo )

ganeshie8 (ganeshie8):

click \(\href{https:///www.google.com/url?sa=t&rct=j&q=&esrc=s&source=web&cd=20&cad=rja&uact=8&ved=0CFQQFjAJOAo&url=http%3A%2F%2Fwww.math.muni.cz%2F~bulik%2Fvyuka%2Fpen-20070711.pdf&ei=suSIVZ-ZL4G9sAWMh7QY&usg=AFQjCNESVFKertjKo1RRDenoVvOsgsul0w}{this}\)

OpenStudy (freckles):

By the way I got my problem from here if you wanted to know : https://www.google.com/url?sa=t&rct=j&q=&esrc=s&source=web&cd=20&cad=rja&uact=8&ved=0CFQQFjAJOAo&url=http%3A%2F%2Fwww.math.muni.cz%2F~bulik%2Fvyuka%2Fpen-20070711.pdf&ei=suSIVZ-ZL4G9sAWMh7QY&usg=AFQjCNESVFKertjKo1RRDenoVvOsgsul0w (by the way do you know if there is a way to hide how long this link actually is using some kind of LaTeX voodoo )

OpenStudy (usukidoll):

what about bit.ly sites?

OpenStudy (freckles):

omg too cute and it was A17 just in case you were scrolling for the probem

OpenStudy (freckles):

@UsukiDoll what is that "bit.ly sites"?

OpenStudy (freckles):

or how do you use it

ganeshie8 (ganeshie8):

@freckles change http:// to http:/// 3 forward slashes are needed

OpenStudy (usukidoll):

bit.ly <--- online site used to shorten long url links

OpenStudy (freckles):

\[\href{ http:///www.wolframalpha.com/input/?i=integrate%28kitty%2Ckitty%29 }{WOLFRAM!} \]

ganeshie8 (ganeshie8):

congrats!!

OpenStudy (freckles):

you can teach an old dog new tricks so congrats to you on doing just that :p

OpenStudy (freckles):

http://bit.ly/1SFjO6a this is the thing @UsukiDoll was talking about

OpenStudy (freckles):

That is kind of neat too

OpenStudy (freckles):

though I kind of like writing my links in cute words

OpenStudy (usukidoll):

bit.ly <--- online site used to shorten long url links

OpenStudy (freckles):

hmmm anyways a mod 3 =0,1, or 2 1 mod 3=1 2 mod 3=2 3 mod 3=0 4 mod 3=1 5 mod 3=2 6 mod 3=0 ... 1,4,...,1+3i mod 3 will give us 1 2,5,...,2+3i mod 3 will give us 2 3,6,...,3+3i mod 3 will give us 3 (which this makes sense) but any of 1+3i,2+3i,3+3i can be even or odd so I don't understand if we assume m is even why we can do mod 3 like what does it tell us doing that